OpenStudy (anonymous):

simplify ((1/5+h)-(1/5))/h

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ 5+h }-\frac{ 1 }{ 5 } }{ h }\] like this?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."